OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points? (1, 7), (10, 1)

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = ( 1- 7)/(10-1) m = -6/9 m = -2/3 So the slope is -2/3

OpenStudy (anonymous):

Thank you!!!

jimthompson5910 (jim_thompson5910):

yw